OpenStudy (anonymous):

what sum of money invested at 9% interest per year is needed to provide an income of £45 per year?

OpenStudy (anonymous):

the principal is compounded annualy ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I mean, the interest is paid once a year right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with 9% p.a.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x * 0.09 = £45

OpenStudy (anonymous):

x * 0.09 = £45 | :0.09 x = £45/0.09 x = £500

OpenStudy (anonymous):

why would you multiply?